Given rms speed, c = (3 RT/M)^1/2 Determine the t
Answer #1

a)

Find T and P in which 3.011*10^23 molecules mass of 6.64*10^-26 kg in a V = 5*10^-3 m3 container and V = 560 m/s

First, let us find T from the equation

RMS = sqrt(3RT/MW)

560 m/s = sqrt(3*8.314 J/molK * T / (MW))

find MW

MW = mass / mol = number of molecules * mass of molecules / mol of molecules

MW = ( 6.64*10^-26)(3.011*10^23) kg /

MW = 0.0199 kg / (0.4349 mol) = 0.045757 kg/mol

560 m/s = sqrt(3*8.314 J/molK * T / (0.045757))

Solve for T

T = (560^2)/(3*8.314)*0.045757 = 575.310 K

T = 575.310-273 = 302.31 °C

For Pressure

apply ideal gas law

PV = nRT

P = nRT/V

n = 0.4349 mol, R = 8.314 J/molK

V = 5*10^-3 m3

P = (0.4349 mol )(8.314 J/molK)(575.310 K)/(5*10^-3 m3) = 416036.41 Pa

1 atm = 101325 Pa

so

P= 4.10 atm

For all further questions, please consider posting them in different set of Q&A. We are only allowed to answer to 1 question per set.